OpenStudy (anonymous):

Find the difference in a altitude between a mountain that has an altitude of 2632 feet and a desert valley that is 143 feet below sea level.

OpenStudy (anonymous):

As sea level is defined as altitude 0, we have: 2632 - (-143) = 2632 + 143 = ?

OpenStudy (anonymous):

2,775

OpenStudy (anonymous):

Hello